OpenStudy (anonymous):

Find an exact value. sine of nineteen pi divided by twelve. quantity square root of two minus square root of six divided by four. quantity square root of six minus square root of two divided by four. quantity square root of six plus square root of two divided by four. quantity negative square root of six minus square root of two divided by four. @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

wut 0_o wuts going on here

OpenStudy (anonymous):

i need help with this

zepdrix (zepdrix):

\[\Large\rm \sin\left(\frac{19\pi}{12}\right)=\sin\left(\frac{19\pi/6}{2}\right)\]We can start there, ya? And maybe apply half-angle formula? What do you think? :3

OpenStudy (anonymous):

okay yes

zepdrix (zepdrix):

\[\Large\rm \sin\left(\frac{\color{orangered}{\theta}}{2}\right)=\pm\sqrt{\frac{1-\cos\color{orangered}{\theta}}{2}}\]

zepdrix (zepdrix):

So then for our problem,\[\Large\rm \sin\left(\frac{\color{orangered}{19\pi/6}}{2}\right)=\pm\sqrt{\frac{1-\cos\color{orangered}{\frac{19\pi}{6}}}{2}}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

What's cosine of 19pi/6? Hmm that angle is too large.. it's larger than 2pi, so we should try to unwind it by subtracting some 2pi's from it

zepdrix (zepdrix):

\[\Large\rm \frac{19\pi}{6}-2\pi\quad=\quad \frac{19\pi}{6}-\frac{12\pi}{6}=\frac{7\pi}{6}\]Ya?

zepdrix (zepdrix):

So our problem becomes,\[\Large\rm \pm\sqrt{\frac{1-\cos\color{orangered}{\frac{7\pi}{6}}}{2}}\]

zepdrix (zepdrix):

All we did was spin around the circle once and land in the same spot, to get a better angle to work with.

zepdrix (zepdrix):

Now simplify +_+

OpenStudy (anonymous):

i have trouble simplifying that..

zepdrix (zepdrix):

bust out your unit circle or something.. to get the cosine value :U requires some memorization to get those down

zepdrix (zepdrix):

cosine is the x coordinate

OpenStudy (anonymous):

\[-\sqrt{6}-\sqrt{2}/4?\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

@zepdrix